OpenStudy (anonymous):

Solve the system of equations: what is the x-value for the solution? 2x + 7y = 3 and x = 1 - 4y -1 5 4 3

OpenStudy (anonymous):

x=5 &y=-1

OpenStudy (jhannybean):

use your second equation to make a substitution for the first equation eq 1 : 2x + 7y = 3 eq 2 : x= 1-4y replace x value in equation 1. 2(1 - 4y) + 7y = 3 distribute. 2(1) - 2(4y) + 7y = 3 2 - 8y + 7y = 3 2 - y = 3 -y = 1 y = -1 now that we've found y, we can plug y back into any of our equations to solve for x. You will then have both of your values.

OpenStudy (anonymous):

thxxxx :) :D